OpenStudy (kkern00):

need serious math help Find the exact value of the following expression cos ^-1 sqrt 3/2

OpenStudy (kkern00):

\[\cos^{-1} \sqrt{3/2}\]

ganeshie8 (ganeshie8):

I think you just need to memorize this value

OpenStudy (kkern00):

theres a certain answer being looked for i just have absolutely no idea how to find it

ganeshie8 (ganeshie8):

cos(30) = sqrt(3)/2

OpenStudy (legomyego180):

He's right, this is something you just have to memorize. https://www.youtube.com/watch?v=LE6dmczMc68

OpenStudy (kkern00):

i'm not even quite sure what this means. I have not taken trig i'm just trying to maybe get explanations on how to solve some problems, but i'll try it out thank you!

OpenStudy (legomyego180):

It's a trigonometric value, the easiest way to know it is to have it memorized. Its like knowing how to calculate the area of a circle or perimeter of a rectangle. There are other ways to figure it out but the easiest way and first way you're usually taught is to just memorize that pi*r^2=area and 2l+2w=perimeter, just like cos(30)=(sqrt(3)/2) sin(30)=(1/2) etc. The cos^-1 just means the inverse, so cos(x)=(sqrt(3)/2) which is 30.

OpenStudy (legomyego180):

Hope that analogy didnt confuse you. Just memorize the first quadrant of the unit circle. It's only a few numbers, not that difficult and youll have to memorize them later anyway.

OpenStudy (kkern00):

i'll work on it thank you (:

OpenStudy (robtobey2):

\[\cos ^{-1}\left(\frac{\sqrt{3}}{2}\right)=\frac{\pi }{6}=60 {}^{\circ} \]

OpenStudy (robtobey2):

From Mathematica:\[\frac{180 \cos ^{-1}\left(\sqrt{\frac{3}{2}}\right)}{\pi }=0.\, +37.7281 i \]

OpenStudy (marcelie):